OpenStudy (anonymous):

If a/x-b/y=c and xy= 1/c, then bx= A) 1-ay B) ay C) ay+1 D) ay-1or E) 2ay ?

OpenStudy (anonymous):

D) ay-1

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Add the first 2 terms to get an xy on the bottom and sub in 1/c for it.

OpenStudy (anonymous):

That's what i did but i get B as the answer which is not correct.

OpenStudy (anonymous):

U did it wrong, u should get (ay - bx) /xy = c as first step.

OpenStudy (anonymous):

This is how I did it: If xy = 1/c, then c = 1/xy If a/x - b/y = c, a/x - b/y = 1/xy Multiply a/x - b/y = 1/xy, by it's LCD which is xy This gives you ay - bx = 1 Add bx to both sides, subtract 1 from both sides. ay -1 = bx or bx = ay -1

OpenStudy (anonymous):

Oh ok tysm for explaining.

OpenStudy (anonymous):

(ay - bx) /xy = c -> (ay - bx) = c * 1/c = 1 -> bx = 1 - ay